OpenStudy (alexishope47):

PLEASE HELP! For some reason this question is stumping me. Solve the following system of equations and show all work. y=x^2 + 3 y=x+5

OpenStudy (alexishope47):

Will Fan and Medal!!

OpenStudy (gherbowife):

go to mathway.com i promise you'll thank me right! :)

OpenStudy (alexishope47):

I tried math way. It great for getting the answer but it doesn't give me an explanation of how it got the answer. I need to show work. :/

OpenStudy (tacotime):

y=x^2 + 3 y=x+5 x^2 + 3 = x + 5 x^2 - x - 2 = 0 now factor: What 2 numbers add to -1 and multiply to -2?

OpenStudy (alexishope47):

1 and -2?

OpenStudy (alexishope47):

So the solutions to the system are x= 1 and x= -2?

OpenStudy (study312):

well if he said which numbers are added together to get -1 then its -3 + 2 = -1 and 1 x -2 = -2. THIS IS A GUESS

OpenStudy (alexishope47):

Well, 1+(-2) is -1 and 1•-2 is -2.

OpenStudy (tacotime):

No @alexishope47 you were right. The answers are x = 1 and x = -2. idk if you know about factoring, but you usually only get two solutions.